OpenStudy (anonymous):

double integral of e^(x+y)/(x-y) dx dy Evaluate the integral where R is the trapezoidal region with corners at (1,0), (2,0), (0,2), (0,-1)

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\int\limits_{?}^{?}e ^{(x+y)/(x-y)}dxdy\]

OpenStudy (anonymous):

Are you sure you typed the corner points right?

OpenStudy (anonymous):

(1,0), (2,0), (0,-2), (0,-1)

OpenStudy (anonymous):

sorry about that.

OpenStudy (anonymous):

You have to divide your integral into two. One from x=0 to x=1 and one from x=1 to x=2

OpenStudy (anonymous):

i don't follow...2 double integrals??

OpenStudy (anonymous):

\[ \int_0^1\int_{x-2}^{x-1} f(x,y) dy dx + \int_1^2\int_{x-2}^{0} f(x,y) dy dx \]

OpenStudy (anonymous):

i don't really understand why you did that but i can kinda see where you are coming from. i don't really understand your bounds

OpenStudy (anonymous):

Wrtie the equations of the two parallel lines of the trapeziods and tell me what they are?

OpenStudy (anonymous):

on the first part of the integration why does the x only go from 0 to 1

OpenStudy (anonymous):

Answer my question above first. It will help you understand

OpenStudy (anonymous):

y=x-2 and y=x-1?

OpenStudy (anonymous):

actually i don't really understand the equations that you got. i would really apprecciate if you could explain it to me

OpenStudy (anonymous):

oh nevermind...got it

OpenStudy (anonymous):

after thinking about it, it would be lot easier to compute your integral after a change of variables. u= x - y and v = x

OpenStudy (anonymous):

that doesn't make sense. then it would be v+y/ u??

OpenStudy (anonymous):

You will have one itntegral to worry about \[ 1\le u \le 2\\ 0\le v \le u \] In the uv plane. Do not forget the jacobian determinant. Your integral will be \[ \frac{3 \left(e^2-1\right)}{4 e} \]

OpenStudy (anonymous):

Yes it does make sense x+y = 2 v -u and x-y =u

OpenStudy (anonymous):

Your integral becomes \[ \int _1^2\int _0^ue^{\frac{2 v}{u}-1}dvdu \]

OpenStudy (anonymous):

The inner integral \[ \int_0^u e^{\frac{2 v}{u}-1} \, dv=\frac{\left(e^2-1\right) u}{2 e} \]

OpenStudy (anonymous):

Integrate what you got above \[ \frac{\left(e^2-1\right) \int_1^2 u \, du}{2 e}=\frac{3 \left(e^2-1\right)}{4 e} \]